OpenStudy (anonymous):

how would i differentiate this? y=3.9657(0.9982^x)

OpenStudy (anonymous):

d(a^x)=a^x*lna dx

OpenStudy (anonymous):

would it be y'=3.9657(0.9982^x) ln0.9982 ?

OpenStudy (anonymous):

\[\frac{d}{dx}a^x=a^x \times \log_{e}a=a^x \times \ln a\]

OpenStudy (anonymous):

do you know why it's like that??

OpenStudy (anonymous):

yes, you can arrive at the formula by applying log to both the sides

OpenStudy (anonymous):

@Nishant_Garg actually i didn't meant you :)

OpenStudy (anonymous):

yes i do thanks! also to find the instantaneous rate of change (i am given a table of values) would I just sub in the x values?

OpenStudy (anonymous):

Oh my bad

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

cool thanks!